OpenStudy (anonymous):

add parenthesies to the equation below so that x =-5 5x4-6/2=x

OpenStudy (anonymous):

i think its 5x(4-6)/2=x

OpenStudy (mertsj):

\[\frac{[5\times(4-6)]}{2}\]